OpenStudy (anonymous):

solve for L R=r+400(W-L)/N

OpenStudy (anonymous):

is the little r over the N as well?

OpenStudy (anonymous):

no its seperate

OpenStudy (anonymous):

R = r + 400((W-L)/N R = r + (400W - 400L)/N R = r + 400W/N - 400L/N R - r - 400W/N = -400L/N N*(R - r - 400W/N) = -400L (N*(R - r - 400W/N))/-400 = L

OpenStudy (anonymous):

To elaborate on my answer: R = r + 400((W-L)/N R = r + (400W - 400L)/N (distributed the 400) R = r + 400W/N - 400L/N (divided both terms in the parenthese by N) R - r - 400W/N = -400L/N (move everything but the term with L, to other side) N*(R - r - 400W/N) = -400L (multiply both sides by N) (N*(R - r - 400W/N))/-400 = L (divide both sides by -400)

OpenStudy (anonymous):

Could you go (NR - Nr - 400W)/-400 = L ?

OpenStudy (anonymous):

yeah, i suppose you could distribute the N if you want also. either way :).